OpenStudy (anonymous):

simplify the expression (x^3y^4 over x^-3y^7)^-2 A. y^3 over x^7 B. x^3 over y C. y^6 over x^10 D. x^-6 over y^-5 Please help

OpenStudy (mertsj):

\[(\frac{x^3y^4}{x ^{-3}y^7})^{-2}\]

OpenStudy (mertsj):

Simplify inside the parentheses first.

OpenStudy (anonymous):

not good at this :(

OpenStudy (mertsj):

When dividing, subtract the exponents.

OpenStudy (anonymous):

x^3?

OpenStudy (mertsj):

\[\frac{x^3}{x ^{-3}}=x ^{3-(-3)}=???\]

OpenStudy (mertsj):

Sorry for the typo.

OpenStudy (mertsj):

What is 3-(-3) ?

OpenStudy (anonymous):

6

OpenStudy (mertsj):

\[\frac{x^3}{x ^{-3}}=\]

OpenStudy (anonymous):

x^6?

OpenStudy (mertsj):

Good job. Now what about this:

OpenStudy (mertsj):

\[\frac{y^4}{y^7}=y ^{4-7}=\]

OpenStudy (anonymous):

y^11

OpenStudy (mertsj):

What is 4-7?

OpenStudy (anonymous):

-3

OpenStudy (mertsj):

You have 4 dollars and you spend 7 dollars

OpenStudy (anonymous):

sorry

OpenStudy (mertsj):

So what is this:

OpenStudy (mertsj):

\[\frac{y^4}{y^7}\]

OpenStudy (anonymous):

y^3

OpenStudy (mertsj):

I thought you just said that 4-7 is -3 and now you are saying it is 3. Which is it?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

lmaoooo

OpenStudy (mertsj):

\[\frac{x^3y^4}{x ^{-3}y^7}=\frac{x^3}{x ^{-3}}\times\frac{y^4}{y^7}=x ^{3-(-3)}\times y ^{4-7}=x^6y ^{-3}\]

OpenStudy (anonymous):

thts not a option

OpenStudy (mertsj):

But remember...that fraction was supposed to be raised to the power of -2.

OpenStudy (mertsj):

So we have: \[(x^6y ^{-3})^{-2}\]

OpenStudy (mertsj):

When you raise to a power, you multiply the exponents. So what do you get when you do that?

OpenStudy (anonymous):

idk :(

OpenStudy (mertsj):

Then I guess you need to figure out how to multiply integers.

OpenStudy (anonymous):

.-.

OpenStudy (mertsj):

I'll give you some practice examples: 1. (6)(3) 2. (-6)(3) 3. (6)(-3) 4. (-6)(-3)